1) Calculate σ x-bar if μ= 2.0, σ=1.6, and n = 4. and What is the value of μ x-bar?

2) Write the binomial probability as a normal probability using the continuity correction. P(x>66).
Answer #1

Solution :

Given that,

1)

mean = = 2

standard deviation = = 1.6

n = 4

The sampling distribution of mean and standard deviation is ,

= 2 and

= / n = 1.6 / 4 = 1.6 / 2 = 0.8

2)

Using the continuity correction,

P(x > 66) = P(x > 66 + 0.5) = P(x > 66.5)